OpenStudy (anonymous):

-7x^2 - 23x + 10 = 0 What are the solutions of the quadratic equation?

OpenStudy (amistre64):

try using the quadratic formula

OpenStudy (amistre64):

given: ax^2 + bx + c = 0 x = -b +- sqrt(b^2 - 4ac) ------------------- 2a

OpenStudy (anonymous):

I think it's the first answer

OpenStudy (dangerousjesse):

Solve for x over the real numbers: \[\large -(7 x^2)-23 x+10 = 0 \]Multiply both sides by a term to make the coefficient of x^2 positive. Multiply both sides by -1: \[\large 7 x^2+23 x-10 = 0 \]Using the quadratic formula, solve for x. \[\large x = \frac{-23±\sqrt{23^2-4×7 (-10)}}{2×7} = \frac{-23±\sqrt{529+280}}{14} = \frac{-23±\sqrt{809}}{14}:\]

OpenStudy (anonymous):

\[23\pm \sqrt{809}\]

OpenStudy (anonymous):

over -14

OpenStudy (dangerousjesse):

I got \(\large \frac{-23+\sqrt{809}}{14}\) or \(\large \frac{-23-\sqrt{809}}{14}\)

OpenStudy (anonymous):

hmm okay